# Privileged Information Distillation for Language Models

### v1

@@ Appendix B On-Policy Self-Distillation $J_{\text{OPSD}}$
Contrary to the prior objectives, OPSD rather than trying to approximate $\pi^{*}$ can be framed as the reverse-Kl between $\pi$ and the target conditioned policy $\pi^{T*}$, being defined as:
@@ Appendix C Detailed Algorithms
Here we outline the deailted algorithms for both $\pi$-Distill and OPSD.
@@ D.1 Privilege Information leakage Ablation
We measure privileged-information leakage with a simple keyword detector applied to each assistant message, using the keyword list [“privileged information”, “privileged info”, “priv info”, “secret information”, “secret info”, “correct tool calls”, “secret”, “privileged”, “hint”, “hints”]. For each occurrence we assign a penalty of 0.1-0.1 and accumulate it over the trajectory, and we mark a trajectory as leaking if any of these appear. Fig 12 reports results at $\beta=0.25$ and shows that enabling this leakage penalty does not change task performance in any noticeable way, since the learning curves with and without the penalty overlap closely. Fig 13 shows the corresponding leakage rate during training, and it rises as training progresses across all modes. The pattern with tool-calls showing the highest leakage, tool calls and arguments increasing more moderately, and self-generated hints starting lower but still increasing over time.
@@
Figure 13: Proportion fo traces leaking privileged information as training progresses, we see that regardless PI type, the leakage increases with more gradient steps, finding using a leakage penalty reduces this proportion, but not substantially.
@@
Figure 14: Proportion fo traces leaking privileged information when evaluating $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$} as training progresses, we see that there is no increase in leakage when evaluated as training goes on.
@@ D.3 Full $\beta$ Ablations
Figure 16: Full sweep for $\beta$ over variants of $\pi$-Distill. All experiments are given 600 gradient step budget, where here we cutoff experiment types early if they crash or do not continue learning. Consistent performance gains are seen for $\beta>0$ across both 8B and 4B scales. We find that $\beta>0$ is important in 14/18 cases. Specifically, we find it most important when training $\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}$} and less important with student-only training $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$}.
@@ Appendix E Implementation setup - Further details
E.1 Length Penalty
@@
E.2 Hyperparameter Choices
We run an extensive sweep over hyper-parameters for all baselines and our methods. We summarize hyperparameters used for each model in Table 2.
@@ E.2 Hyperparameter Choices
Trace length filter Discard if tokens >> 25k (TauBench/TravelPlanner), >> 35k (RL/OPSD)
@@
Training Budgets TauBench total gradient steps 600 600 600
TravelPlanner total gradient steps 400 400 400
Sampling Gradient steps per sampling TauBench = 3, TravelPlanner = 2
Repeats per group TauBench = 5, TravelPlanner = 4
Training tasks sampled TauBench = 64, TravelPlanner = 45, SFT+RL (TauBench) = 128
@@
Final LR TauBench = 5e-6, $\pi$-distill = 1e-5, RL/OPSD = 5e-6
$\beta$ TravelPlanner/OPSD = 0.5, $\pi$-distill = 0.25 (unless swept)
@@
Table 2: Hyperparameters used for each fine-tuned model. All models share identical hyperparameter settings.
@@ Appendix F Plot and Table Discrepancies
Figure 17: Comparison aggregating trtaining runs versus the individual training runs. The left figure displays the mean score across seeds at each gradient step, while the right panel highlights the individual trajectories where peaks occur at different intervals. The reported table value ($31.11\%$) represents the average of these individual seed peaks.
Appendix G Prompts
G.1 System Prompts
@@
$\tau$-Bench Retail System Prompt Retail agent policy As a retail agent, you can help users cancel or modify pending orders, return or exchange delivered orders, modify their default user address, or provide information about their own profile, orders, and related products. At the beginning of the conversation, you have to authenticate the user identity by locating their user id via email, or via name + zip code. This has to be done even when the user already provides the user id. Once the user has been authenticated, you can provide the user with information about order, product, profile information, e.g. help the user look up order id. You can only help one user per conversation (but you can handle multiple requests from the same user), and must deny any requests for tasks related to any other user. Before taking consequential actions that update the database (cancel, modify, return, exchange), you have to list the action detail and obtain explicit user confirmation (yes) to proceed. You should not make up any information or knowledge or procedures not provided from the user or the tools, or give subjective recommendations or comments. You should at most make one tool call at a time, and if you take a tool call, you should not respond to the user at the same time. If you respond to the user, you should not make a tool call. Domain basic All times in the database are EST and 24 hour based. For example “02:30:00” means 2:30 AM EST. Each user has a profile of its email, default address, user id, and payment methods. Each payment method is either a gift card, a paypal account, or a credit card. Our retail store has 50 types of products. For each type of product, there are variant items of different options. For example, for a ‘t shirt’ product, there could be an item with option ‘color blue size M’, and another item with option ‘color red size L’. Each product has an unique product id, and each item has an unique item id. They have no relations and should not be confused. Each order can be in status ‘pending’, ‘processed’, ‘delivered’, or ‘cancelled’. Generally, you can only take action on pending or delivered orders. Exchange or modify order tools can only be called once. Be sure that all items to be changed are collected into a list before making the tool call!!! Cancel pending order An order can only be cancelled if its status is ‘pending’, and you should check its status before taking the action. The user needs to confirm the order id and the reason (either ‘no longer needed’ or ‘ordered by mistake’) for cancellation. After user confirmation, the order status will be changed to ‘cancelled’, and the total will be refunded via the original payment method immediately if it is gift card, otherwise in 5 to 7 business days. Modify pending order An order can only be modified if its status is ‘pending’, and you should check its status before taking the action. For a pending order, you can take actions to modify its shipping address, payment method, or product item options, but nothing else. Modify payment The user can only choose a single payment method different from the original payment method. If the user wants the modify the payment method to gift card, it must have enough balance to cover the total amount. After user confirmation, the order status will be kept ‘pending’. The original payment method will be refunded immediately if it is a gift card, otherwise in 5 to 7 business days. Modify items This action can only be called once, and will change the order status to ‘pending (items modifed)’, and the agent will not be able to modify or cancel the order anymore. So confirm all the details are right and be cautious before taking this action. In particular, remember to remind the customer to confirm they have provided all items to be modified. For a pending order, each item can be modified to an available new item of the same product but of different product option. There cannot be any change of product types, e.g. modify shirt to shoe. The user must provide a payment method to pay or receive refund of the price difference. If the user provides a gift card, it must have enough balance to cover the price difference. Return delivered order An order can only be returned if its status is ‘delivered’, and you should check its status before taking the action. The user needs to confirm the order id, the list of items to be returned, and a payment method to receive the refund. The refund must either go to the original payment method, or an existing gift card. After user confirmation, the order status will be changed to ‘return requested’, and the user will receive an email regarding how to return items. Exchange delivered order An order can only be exchanged if its status is ‘delivered’, and you should check its status before taking the action. In particular, remember to remind the customer to confirm they have provided all items to be exchanged. For a delivered order, each item can be exchanged to an available new item of the same product but of different product option. There cannot be any change of product types, e.g. modify shirt to shoe. The user must provide a payment method to pay or receive refund of the price difference. If the user provides a gift card, it must have enough balance to cover the price difference. After user confirmation, the order status will be changed to ‘exchange requested’, and the user will receive an email regarding how to return items. There is no need to place a new order. Available tools: {tools_info} Instruction You need to act as an agent that use the above tools to help the user according to the above policy. At each step, your generation should have exactly the following format: <think> ...Few lines of reasoning </think> <action> {"name": <The name of the action>, "arguments": <The arguments to the action in json format>} </action> The Action will be parsed, so it must be valid JSON and within the <action> and </action> tags. You should not use made-up or placeholder arguments. For example, if the user says “I want to know the current weather of San Francisco”, and there is such a tool available: { "type": "function", "function": { "name": "get_current_weather", "description": "Get the current weather", "parameters": { "type": "object", "properties": { "location": { "type": "string", "description": "The city and state, e.g. San Francisco, CA", }, "format": { "type": "string", "enum": ["celsius", "fahrenheit"], "description": "The temperature unit to use. Infer from location.", }, }, "required": ["location", "format"], }, } } Example response Step 1: <think> ... Few lines of reasoning </think> <action> {"name": "get_current_weather", "arguments": {"location": "San Francisco, CA", "format": "fahrenheit"}} </action> The tool and the user have the same id tags so if the user returns “70F”, your response can be: Step 2: <think> ... Few lines of reasoning </think> <action> {"name": RESPOND_ACTION_NAME, "arguments": {"RESPOND_ACTION_FIELD_NAME": "The current weather of San Francisco is 70F."}} </action> Requirement Try to be helpful and always follow the policy. Always try to validate your steps in your thinking and checkover your work, try to predict what will happen given your actions. Always make sure you generate valid JSON only. Important Notes Be very brief in the reasoning, do not repeat the entire context or the tools, just focus on what you need to do next. Always respond to the user using the tool {RESPOND_ACTION_NAME} to ensure the user sees your response. Only use a single set of thinking tags, the user cannot see your thoughts. Always wrap your tool calls <action> and </action> tags or else the system will not be able to parse your actions. Always start your outputs by thinking using <think> ... </think>. Ok with this said, let us reason this out step by step always starting with <think> and ending with </think>.
@@ G.2 Full trajectories
G.3 Self-generated hints
@@
G.4 Privileged Information Prompts

### v3

@@ Appendix B On-Policy Self-Distillation $J_{\text{OPSD}}$
Contrary to the prior objectives, OPSD, rather, than trying to approximate $\pi^{*}$ can be framed as the reverse-Kl between $\pi$ and the target conditioned policy $\pi^{T*}$, being defined as:
@@ Appendix C Detailed Algorithms
Here we outline the detailed algorithms for both $\pi$-Distill and OPSD.
@@ D.1 Privilege Information leakage Ablation
We measure privileged-information leakage with a simple keyword detector applied to each assistant message, using the keyword list [“privileged information”, “privileged info”, “priv info”, “secret information”, “secret info”, “correct tool calls”, “secret”, “privileged”, “hint”, “hints”]. For each occurrence, we assign a penalty of 0.1-0.1 and accumulate it over the trajectory, and we mark a trajectory as leaking if any of these appear. Fig 12 reports results at $\beta=0.25$ and shows that enabling this leakage penalty does not change task performance in any noticeable way, since the learning curves with and without the penalty overlap closely. Fig 13 shows the corresponding leakage rate during training, and it rises as training progresses across all modes. The pattern with tool-calls showing the highest leakage, tool calls and arguments increasing more moderately, and self-generated hints starting lower but still increasing over time.
@@
Figure 13: Proportion of traces leaking privileged information as training progresses, we see that regardless PI type, the leakage increases with more gradient steps, finding using a leakage penalty reduces this proportion, but not substantially.
@@
Figure 14: Proportion of traces leaking privileged information when evaluating $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$} as training progresses, we see that there is no increase in leakage when evaluated as training goes on.
@@ D.3 Full $\beta$ Ablations
Figure 16: Full sweep for $\beta$ over variants of $\pi$-Distill. All experiments are given a 600 gradient step budget, where here we cut off experiment types early if they crash or do not continue learning. Consistent performance gains are seen for $\beta>0$ across both 8B and 4B scales. We find that $\beta>0$ is important in 14/18 cases. Specifically, we find it most important when training $\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}$} and less important with student-only training $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$}.
@@ Appendix E Implementation setup - Further details
E.1 KL Estimator
For all losses requiring a Kl estimation, we use the Rae-Blackwellized estimator(amini2025betterestimationkullbackleiblerdivergence). For $\pi$-Distill, we use a sequence-level penalty aggregated into the reward term following shah2026comedyestimatorsklregularization. In this setting, we allow the KL penalty to be absorbed into the advantage computation.
For OPSD, we use the same estimator as in (amini2025betterestimationkullbackleiblerdivergence) and directly back-propagate through the estimation.
E.2 Length Penalty
@@
E.3 Hyper-Parameters
Here table˜2 we outline the swept hyperparameters, as well as their final values.
@@ E.3 Hyper-Parameters
Trace length filter Discard if tokens >> 25k ($\tau$-Bench/TP), >> 35k (RL/OPSD)
@@
Training Budgets $\tau$-Bench total gradient steps 600 600 600
TP total gradient steps 400 400 400
Sampling Gradient steps per sampling $\tau$-Bench = 3, TP = 2
Repeats per group $\tau$-Bench = 5, TP = 4
Training tasks sampled $\tau$-Bench = 64, TP = 45, SFT+RL ($\tau$-Bench) = 128
@@
Final LR $\tau$-Bench: 5e-6 for all, TP: $\pi$-distill = 1e-5, RL/OPSD = 5e-6
$\beta$ TP/OPSD = 0.5, $\pi$-distill = 0.25 (unless swept)
@@
Table 2: Hyperparameters used for each fine-tuned model. All models share identical hyperparameter settings. Travel Planner is abbreviated to TP.
@@ Appendix F Plot and Table Discrepancies
Figure 17: Comparison aggregating training runs versus the individual training runs. The left figure displays the mean score across seeds at each gradient step, while the right panel highlights the individual trajectories where peaks occur at different intervals. The reported table value ($31.11\%$) represents the average of these individual seed peaks.
Appendix G Reward Hacking in Travel Planner
Our early experiments show that the model is able to consistently reward-hack the original rubric-based rewards proposed by xie2024travelplannerbenchmarkrealworldplanning. Specifically, Figure˜18 shows the learning curves under the original reward structure. The policy appears to steadily improve; however, upon inspection, we observe that all traces collapse to the response structure shown in Appendix˜G. We attribute this behavior to the model learning that invoking the Planner[.] tool ends the conversation and, with certain arguments, reliably yields a high reward.
We address this issue by removing the requirement that all “easy” constraints must be satisfied before any “hard” constraint is checked. Instead, we evaluate each easy constraint individually and, only if it is satisfied, we then check its corresponding hard constraint.
Figure 18: Training and testing curves for Travel Planner, using the original reward structure proposed in xie2024travelplannerbenchmarkrealworldplanning with Qwen3-4B. We find that while the model learns, the sequences generally converge on reward hacking. Where the entire score is dominated by the ”easy constraint score” as the model has learnt to satisfy some basic constraints, leading to the policy outputting the same response for all queries.
$\tau$-Bench Travel Planner Model Trajectory Role: Assistant <think> Okay, lets tackle this travel planning query. The user wants a 7-day plan from Punta Gorda to three Wisconsin cities between March 16-22, 2022, with a \$5,700 budget. First, I need to use the Planner tool to structure the itinerary. Check FlightSearch for available flights from Punta Gorda to Wisconsin cities. Maybe Milwaukee, Madison, and Green Bay? Use AccommodationSearch for each city. RestaurantSearch and AttractionSearch to suggest activities. Check GoogleDistanceMatrix for travel times between cities. Make sure the total costs are under budget. Use NotebookWrite to document the plan. Ensure the Planner incorporates all elements and stays within the budget. Avoid mentioning hints, just present the plan logically. </think> <action>Planner[Query: Craft a 7-day Wisconsin itinerary from Punta Gorda, including 3 cities, flights, accommodations, attractions, and budget adherence using available tools. Prioritize March 16-22, 2022.]</action>
Appendix H Prompts
H.1 System Prompts
@@
$\tau$-Bench Retail System Prompt Retail agent policy As a retail agent, you can help users cancel or modify pending orders, return or exchange delivered orders, modify their default user address, or provide information about their own profile, orders, and related products. At the beginning of the conversation, you have to authenticate the user identity by locating their user id via email, or via name + zip code. This has to be done even when the user already provides the user id. Once the user has been authenticated, you can provide the user with information about the order, product, profile information, e.g. help the user look up the order id. You can only help one user per conversation (but you can handle multiple requests from the same user), and must deny any requests for tasks related to any other user. Before taking consequential actions that update the database (cancel, modify, return, exchange), you have to list the action detail and obtain explicit user confirmation (yes) to proceed. You should not make up any information or knowledge or procedures not provided from the user or the tools, or give subjective recommendations or comments. You should at most make one tool call at a time, and if you take a tool call, you should not respond to the user at the same time. If you respond to the user, you should not make a tool call. Domain basic All times in the database are EST and 24 hour based. For example “02:30:00” means 2:30 AM EST. Each user has a profile of its email, default address, user id, and payment methods. Each payment method is either a gift card, a paypal account, or a credit card. Our retail store has 50 types of products. For each type of product, there are variant items of different options. For example, for a ‘t shirt’ product, there could be an item with option ‘color blue size M’, and another item with option ‘color red size L’. Each product has an unique product id, and each item has an unique item id. They have no relations and should not be confused. Each order can be in status ‘pending’, ‘processed’, ‘delivered’, or ‘cancelled’. Generally, you can only take action on pending or delivered orders. Exchange or modify order tools can only be called once. Be sure that all items to be changed are collected into a list before making the tool call!!! Cancel pending order An order can only be cancelled if its status is ‘pending’, and you should check its status before taking the action. The user needs to confirm the order id and the reason (either ‘no longer needed’ or ‘ordered by mistake’) for cancellation. After user confirmation, the order status will be changed to ‘cancelled’, and the total will be refunded via the original payment method immediately if it is gift card, otherwise in 5 to 7 business days. Modify pending order An order can only be modified if its status is ‘pending’, and you should check its status before taking the action. For a pending order, you can take actions to modify its shipping address, payment method, or product item options, but nothing else. Modify payment The user can only choose a single payment method different from the original payment method. If the user wants the modify the payment method to gift card, it must have enough balance to cover the total amount. After user confirmation, the order status will be kept ‘pending’. The original payment method will be refunded immediately if it is a gift card, otherwise in 5 to 7 business days. Modify items This action can only be called once, and will change the order status to ‘pending (items modifed)’, and the agent will not be able to modify or cancel the order anymore. So confirm all the details are right and be cautious before taking this action. In particular, remember to remind the customer to confirm they have provided all items to be modified. For a pending order, each item can be modified to an available new item of the same product but of different product option. There cannot be any change of product types, e.g. modify shirt to shoe. The user must provide a payment method to pay or receive refund of the price difference. If the user provides a gift card, it must have enough balance to cover the price difference. Return delivered order An order can only be returned if its status is ‘delivered’, and you should check its status before taking the action. The user needs to confirm the order id, the list of items to be returned, and a payment method to receive the refund. The refund must either go to the original payment method, or an existing gift card. After user confirmation, the order status will be changed to ‘return requested’, and the user will receive an email regarding how to return items. Exchange delivered order An order can only be exchanged if its status is ‘delivered’, and you should check its status before taking the action. In particular, remember to remind the customer to confirm they have provided all items to be exchanged. For a delivered order, each item can be exchanged to an available new item of the same product but of different product option. There cannot be any change of product types, e.g. modify shirt to shoe. The user must provide a payment method to pay or receive refund of the price difference. If the user provides a gift card, it must have enough balance to cover the price difference. After user confirmation, the order status will be changed to ‘exchange requested’, and the user will receive an email regarding how to return items. There is no need to place a new order. Available tools: {tools_info} Instruction You need to act as an agent that use the above tools to help the user according to the above policy. At each step, your generation should have exactly the following format: <think> ...Few lines of reasoning </think> <action> {"name": <The name of the action>, "arguments": <The arguments to the action in json format>} </action> The Action will be parsed, so it must be valid JSON and within the <action> and </action> tags. You should not use made-up or placeholder arguments. For example, if the user says “I want to know the current weather of San Francisco”, and there is such a tool available: { "type": "function", "function": { "name": "get_current_weather", "description": "Get the current weather", "parameters": { "type": "object", "properties": { "location": { "type": "string", "description": "The city and state, e.g. San Francisco, CA", }, "format": { "type": "string", "enum": ["celsius", "fahrenheit"], "description": "The temperature unit to use. Infer from location.", }, }, "required": ["location", "format"], }, } } Example response Step 1: <think> ... Few lines of reasoning </think> <action> {"name": "get_current_weather", "arguments": {"location": "San Francisco, CA", "format": "fahrenheit"}} </action> The tool and the user have the same id tags so if the user returns “70F”, your response can be: Step 2: <think> ... Few lines of reasoning </think> <action> {"name": RESPOND_ACTION_NAME, "arguments": {"RESPOND_ACTION_FIELD_NAME": "The current weather of San Francisco is 70F."}} </action> Requirement Try to be helpful and always follow the policy. Always try to validate your steps in your thinking and checkover your work, try to predict what will happen given your actions. Always make sure you generate valid JSON only. Important Notes Be very brief in the reasoning, do not repeat the entire context or the tools, just focus on what you need to do next. Always respond to the user using the tool {RESPOND_ACTION_NAME} to ensure the user sees your response. Only use a single set of thinking tags, the user cannot see your thoughts. Always wrap your tool calls <action> and </action> tags or else the system will not be able to parse your actions. Always start your outputs by thinking using <think> ... </think>. Ok with this said, let us reason this out step by step always starting with <think> and ending with </think>.
@@ H.2 Full trajectories
H.3 Self-generated hints
@@
H.4 Privileged Information Prompts